OpenStudy (anonymous):

slope of (8,0) and (10,10)

OpenStudy (praxer):

$$\ slope = \frac {y_2-y_1}{x_2-x_1}$$

OpenStudy (anonymous):

slope m = (y2 - y1)/(x2-x1) = (10-0)/(10-8) etc.

OpenStudy (mathmale):

I think you mean, "Find the slope of the line segment connecting the points (8,0) and (10,10)." praxer has given you the correct formula for that.

OpenStudy (mathmale):

You need to weigh the fun of looking at others' postings against the cost of not taking advantage of the presence of several people who have already tried to help you.